OpenStudy (anonymous):

Deon has 12 milliliters of 5% solution. He also has a solution that has a concentration of 30%. How many milliliters of the 30% solution does Deon need to add to the 5% solution to obtain a 20% solution

OpenStudy (anonymous):

if he adds both solutions, the "average conentration" should be 20% average: a% + b% /2 we have a "weighted average", instead of just counting "a" and "b" once, and divide by two, we count the exact millilitres, and divide by total ml weighted average: x * 5% +y*30% = 20% /(x+y) if x and y were =1, then it would be the usual average. instead they are now weighted.

OpenStudy (anonymous):

we can solve the equation by letting the 5% solution =12 ml and re-arranging the terms: 12 * 5 +y*30 = 20 /(12+y) | *(12+y) 60 (12+y) + 30y (12+y) = 20 720 + 60y + 360 y + 30y² = 20 30 y ² + 380 y + 720 = 20 <--- we can set that to =0, and use pq

OpenStudy (anonymous):

30 y ² + 380 y + 700 = 0 change to standard form for quadratic formula y^2 + 12.67 y + 23.33 b=12.67 c=23.33 \[x = -\frac{-b\pm\sqrt{b^2-4c}}{2a} \]\[x = -\frac{-12.67\pm\sqrt{12.67^2-4\times 23.33}}{2} \] \[x_1 = -\frac{-12.67\pm\sqrt{67}}{2} \]\[x_2 = -\frac{-12.67\pm\sqrt{253}}{2} \]

OpenStudy (anonymous):

\[x_1 = -\frac{-12.67+\sqrt{160}}{2} = \frac{-12.67+12.64}{2} \approx0\]\[x_2 = -\frac{-12.67-\sqrt{160}}{2} = -\frac{-12.67-12.64}{2}=-12.65.\]yeah well, that can't be right. the ml of the second solution has to be positive

OpenStudy (anonymous):

Don't you have to do a table

OpenStudy (anonymous):

can you help me do a table

OpenStudy (anonymous):

the formula should be: 60 +30y = 20 /(12+y) I'm not sure where to start...

OpenStudy (anonymous):

maybe we should start with second solution = 0 ml ? :)

OpenStudy (anonymous):

function: concentration =(x*5 + 30y) /(x+y) with x=12 ml = (60 + 30y) /(12+y) second solution (ml) | concentration level (%) ----------------------------------------- 0 | 5% 1 | 6.92% 2 | 8.57% 3 | 10% 4 | 11.25% 12 | 17.5% 13 | 18.0% 14 | 15 | 18.88% 16 | 19.28% 17 | 19.65% 18 | 20.00% <------------------- 19 | 20.32%

OpenStudy (anonymous):

it's also easy to find the result using algebra, if you do it right. my algebra above was broken. correct way: \[\frac{12\times 5\% +y\times 30\% }{12+y} =20\% \]then, multiply both sides with 12+y (I had it divided, that was the error) \[60 +30y =20 (12+y) \]this can be simplified to: \[y=18 ml.\]